# Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview

### v2

@@ 5.2 Nonstationarity of STBNs
Nonstationarity is a significant topic in researches on time series. Actually, however, the reasons inducing it are likely to be different from the view of causality. Three of the most currently popular reasons are:
@@
    Trending: The time series have ascending or descending trending for long terms.
@@
    Time-varying Variance: The variances of additive noise are time-varying, that is also called as heterogeneous causality in papers [89, 90].
@@
Figure 10: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed, thus, they are observable, instead of hidden variable.
@@
For the trending time series, an effective approach is to calculate differences of time series in preprocessing. This approach enables to obtain the differential time series, that is the observations of system dynamics, like Eq. (14). This implied, the dynamics is stable during the period of time, and the structures of STBNs do not change, or in another word, the differential time series are stationary. Thus, if the dynamics equations, e.g., Eq. (14), are not equal to zero, the causal links in STBNs are identifiable through the Jacobian matrix of dynamics [91, 92, 93, 94].
@@
For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [88]. That is, if there is no causal connection between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90]. As shown in Fig. 10, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 10. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and t1t-1, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, it is causally identifiable. Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [90].
@@
Nonstationary causality is challenging but critical, especially for the case of frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 10. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify a invariant structure during this period. However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections. Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [101, 102]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
@@ 5.3 STBNs and Complex networks
Here, we suppose STBNs model has been built, and it is associated with interactive time-series process of multivariate (or multiple units). Then, we further introduce complex network, another type of network model, to view the causal interactions in complex system. The network dynamics for variables $X_{i},\dots,X_{n}$ can be defined as
@@
Figure 11: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
Thus, the first-order STBNs have some good properties if network dynamics have. For example, if we want to know how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [106, 107] or the scale-free property [108] of complex networks. Moreover, the counterfactual outcomes would not change because network systems have tolerance from external “interventions” [109, 110]. We would also know the synchronization [111, 112, 113, 114], controllability [115, 116, 117, 118, 119, 120, 121, 122], resilience [23, 123] of STBNs. We would also be able to sparsely identify the network structure even if the network size is large [24, 25], because the independence are equivalent to predictability in STBNs [21]. Also, the network size of STBNs can be reduced, because many complex networks can be simplified and still provide an insightful description of the causality of interest [124, 125]. This is meaningful to design counterfactual inference algorithm on smaller networks, and then obtain higher inference efficiency.
@@ 6 Conclusion
In this work, we mainly focus on the spatio-temporal graphical counterfactuals, and organize an overview for it to discuss its theoretical foundations and application approaches. To discuss the theoretical foundations, a survey is investigated, and the definition of counterfactuals is defined based on the concept of potential outcome in the framework of POM firstly. And then, the framework of SCM is also introduced to infer counterfactuals through graphical languages, which are equivalent to the counterfactuals in POM. Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it is able to recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$. With this algorithm, the spatio-temporal graphical counterfactuals can be inferred if have two elements, that is, the network structure of STBNs and the corresponding algorithm to identify the network structure. This work not only discusses the identification algorithms in various nonstationary cases, but also discusses the feasibility to improve the algorithm efficiency from the view of complex networks.
@@
This work is supported by the National Key R&D Program of China under Grant No. 2022ZD0120003, the National Natural Science Foundation of China under Grant Nos. 62233004, 62273090, 52525204, 52572354, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002, the US National Science Foundation under Grant No. 2047488 and the Rensselaer-IBM AI Research Collaboration.

### v3

@@ 5.2 Nonstationarity of STBNs
Nonstationarity is a significant topic in researches on time series. However, the reasons for it are likely to be diff from the view of causality. Three of the most popular reasons are:
@@
    Trending: The time series has ascending or descending trending for long terms.
@@
    Time-varying Variance: The variances of the additive noise are time-varying, a phenomenon often referred to as heterogeneous causality in the literature [89, 90].
@@
Figure 10: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed; Thus, they are observable, instead of a hidden variable.
@@
For trending time series, an effective approach is to calculate the differences of the time series during preprocessing. This approach obtains the differential time series, that is, the observations of the system dynamics, as in Eq. (14). This implies that the dynamics is stable over the period, and that the structures of STBNs do not change, that is, the differential time series are stationary. Thus, if the dynamics equations, e.g., Eq. (14), are not equal to zero, the causal links in STBNs are identifiable through the Jacobian matrix of dynamics [91, 92, 93, 94].
@@
In the case of time-varying variance, causal identifiability relies on the mutual independence of exogenous additive noises [88]. Specially, if no causal connection exists between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, treating the timestamps at each time step as observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90]. As shown in Fig. 10, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 10. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and t1t-1, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, causal identifiability is preserved. Note that, however, the premise to do so is that any hidden confounders associated with time-varying variance can be expressed as smooth or non-smooth functions of timestamps [90].
@@
Nonstationary causality presents a significant challenge, particularly in scenarios involving frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 10. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify an invariant structure during this period. However, if network changes are frequent, the samples used for identification over a short period would be few. In contrast, we usually need sufficient samples to assess the significance of causal connections. Thus, another reasonable approach is to first identify the change points, and then, determine the causality for the corresponding periods [101, 102]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are other approaches that view nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
@@ 5.3 STBNs and Complex networks
Here, we suppose STBNs model has been built, and it is associated with an interactive time-series process of multivariate (or multiple units). Then, we introduce a complex network, another network model, to examine the causal interactions in a complex system. The network dynamics for variables $X_{i},\dots,X_{n}$ can be defined as
@@
Figure 11: Diagram of discretized network dynamics. Here, AA is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
First-order STBNs exhibit advantageous properties when the underlying network dynamics process certain structural features. For example, determining how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [106, 107] or the scale-free property [108] of complex networks. Moreover, counterfactual outcomes would not change because network systems have tolerance from external “interventions” [109, 110]. We would also know the synchronization [111, 112, 113, 114], controllability [115, 116, 117, 118, 119, 120, 121, 122], resilience [23, 123] of STBNs. We would also be able to sparsely identify the network structure even if the network size is large [24, 25], because the independence is equivalent to predictability in STBNs [21]. Also, the network size of STBNs can be reduced, because many complex networks can be simplified and still provide an insightful description of the causality of interest [124, 125]. This reduction supports the design of a counterfactual inference algorithm on smaller networks, thereby improving inference efficiency.
@@ 6 Conclusion
This work focuses on the spatio-temporal graphical counterfactuals and presents an overview of their theoretical foundations and application approaches. To discuss the theoretical basis, a survey is conducted, and the definition of counterfactuals is based on the concept of potential outcomes in the POM framework. We then, introduce the SCM framework, which infers counterfactuals via graphical languages equivalent to those in POM. Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it can recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$. With this algorithm, spatio-temporal graphical counterfactuals can be inferred, given two elements: the network structure of STBNs and the algorithm identifying it. This work discusses structure identification algorithms under various nonstationary conditions and examines the feasibility of improving algorithm efficiency from the perspective of complex networks.
@@
This work was supported in part by the National Key R&D Program of China (Grant No. 2022ZD0120004), the Youth Scientist Project of the Ministry of Science and Technology of China (Grant No. 2025YFF0524100), the National Natural Science Foundation of China (Grant Nos. 62233004, 62273090, 62073076, T2541017, 52525204, 52572354), the Zhishan Youth Scholar Program of the Southeast University, the Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002), the basic research program of Jiangsu (Grant No. BK20253020), the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B54), the US National Science Foundation under (Grant No. 2047488) and the Rensselaer-IBM AI Research Collaboration.